Solve the system using substitution.

y = 2x + 11
y + x = 5

Answer by etutorworld(19)

Solve the system using substitution.

y = 2x + 11
y + x = 5
To find the value of variable we substitute the value of 'y' from 1st equation into the second equation, we get
(2x + 11) + x =5
3x + 11 = 5
3x = 5-11
3x = -6
x = -2
To find the value of y plug x = -2 in any of the given equations
y = 2(-2) + 11
= -4 +11
y = 7
Solution : (x, y) = ( -2 ,7 )